# Susceptibilities of rotating quark matter in Fourier-Bessel basis

### v1

@@ Abstract
preprint: RIKEN-iTHEMS-Report-25
I Introduction
@@
Rotating QCD matter has garnered significant attention in recent years, particularly in the context of heavy-ion collisions, which create the most vortical fluid known in nature [1]. On the theoretical side, the rigidly rotating system is an idealized setup in order to understand fundamental features of rotating relativistic matter [2, 3, 4], and has already revealed typical rotational responses of QCD matter in terms of the chiral symmetry breaking [5, 6, 7, 8, 9, 10, 11, 12], the deconfinement transition [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], both of them [24, 25, 26, 27], and other possible transitions [28, 29, 30, 31, 32]; see also the corresponding lattice studies in Ref. [33, 34, 35, 36, 37, 38, 39, 40]. Even under such simplification, however, accessible aspects of rotating QCD matter are still limited. In particular, since the first principle simulation under rotation suffers from a sign problem [41], the sophistication of the model approaches and perturbative computation is an inevitable direction.
@@
A difficulty we face in field-theoretical studies of relativistic rotating matter is the coordinate-dependence of physical observables (see also Refs. [5, 42] for an exceptional case under a strong magnetic field). As demonstrated in preceding works, maintaining causality in a relativistic rotating system necessitates the consideration of a finite-size geometry, typically a cylinder [43]. Such a finite-size system violates translational invariance along the radial direction, leading to coordinate-dependent thermodynamic quantities. In more field-theoretical language, propagators of rotating matter are described not by the plane-wave basis, but the Fourier-Bessel basis. As a result, in an interacting theory under rotation, it is necessary to carefully revisit to standard schemes employed in the usual nonrotating case, including the Feynman rules [44].
@@
A typical modification is found in effective models describing phase transition under rotation. The standard mean-field analysis is not directly applied to radially inhomogeneous order parameters. One way to bypass this issue is to utilize the local density approximation, where the radial modulation of the order parameter is supposed to be negligible. Under this approximation, for instance, the mean-field analysis for the spontaneous chiral symmetry breaking can be performed parallelly to the usual case, but the constant dynamical mass is replaced with the radial-coordinate dependent one [6]. A remarkable feature of this analysis is that the self-consistent gap equation involves only the single radial coordinate (through the one-loop of a single propagator), and this is the reason why such a simple analysis works. However, the validity of this strategy becomes subtle when we argue mesonic fluctuations and susceptibilities, which consist of multi-point propagators, and thus involve more radial coordinates than one.
@@
In order to deal with such an issue, Ward-Takahashi (WT) identities would provide a key computational guide. WT identities are highly useful in investigations of QCD properties at finite temperature and density, where Lorentz symmetry is violated by external conditions. For instance, a WT identity in term of the chiral symmetry relates the quark condensate to the two-point correlation function derived from quark bilinear operators corresponding to the pion mode (referred to as the pion susceptibility), and thus plays a crucial role in assessing the consistency within an effective model approach [45, 46, 47]. Furthermore, a WT identity for the U(1) axial symmetry associates quark two-point correlation functions with the topological susceptibility, which is originally described by gluonic degrees of freedom [48, 49, 50, 45, 46]. Since these WT identities related to the global symmetries is irrelevant to the spacetime geometry, they are expected to be powerful constraints even in the quantum field theory for rotating matter.
@@
The present work aims to analyze fermionic two-point correlation functions of rotating matter in the Fourier-Bessel basis, within the mean-field level of an Nambu-Jona-Lasinio (NJL) model. We compute various quark susceptibilities, including meson, baryon, and topological susceptibilities. In a similar computational manner, we can evaluate the density of moment of inertia in the interacting theory, while the net moment of inertia has already derived in several preceding studies [8, 51]. Besides, we show that a WT identity enforces the mesonic susceptibilities to take the similar form to that in the translational invariant systems. These analyses not only establish the groundwork for more comprehensive treatments of quantum fluctuations in rotating matter, but also provides new benchmarks to characterize the rotational response to QCD phase transition. Indeed, without rotation, the phase structure of hot QCD has been explored, revealing the chiral crossover where the quark condensate decreases smoothly with increasing temperature [52, 53]. Additionally, meson susceptibilities, derived from quark bilinear correlation functions, have also been investigated and have provided valuable insights into meson property fluctuations and symmetry breaking, including chiral $SU(2)$ and $U(1)$ axial symmetries [54, 53, 55, 56, 57]. Given this fact, our study extends the understanding of rotating QCD matter by elucidating how rotation modifies these susceptibility functions and the associated symmetry restoration patterns.
@@
This paper is organized as follows. In section II, we present the theoretical framework in the rotating cylindrical frame. We briefly review the formulation of the Dirac propagator described in the Fourier-Bessel basis and introduce the definition of the susceptibility functions. The boundary condition in a finite cylinder is also provided. In section III, we analyze the susceptibility functions within the NJL model under the mean-field approximation. We examine the implementation of the local density approximation in the current framework and explore the associated constraints on the quark propagator. This approximation simplifies the calculation of the susceptibility functions and, guided by a WT identity, makes it possible to evaluate meson susceptibilities. In section IV, we show our numerical results on the susceptibility functions at finite temperature under rotation. Finally, we summarize our findings and discuss possible future directions in section V.
II Theoretical setup
@@
II.1 Dirac propagator in the Fourier-Bessel basis
@@
$\Bigl{[}i\gamma^{\mu}(\partial_{\mu}+\Gamma_{\mu})-m\Bigr{]}\psi=0,$ (1)
@@
$g_{\mu\nu}=\eta_{ij}e^{i}_{\mu}e^{j}_{\nu}.$ (2)
In this paper, we consider a system rotating around the $z$-axis with an angular velocity, ${\bm{\Omega}}=\Omega\hat{\bm{z}}$, and thus employ the corresponding metric as
$g_{\mu\nu}=\begin{pmatrix}1-(x^{2}+y^{2})\Omega^{2}&y\Omega&-x\Omega&0\\ y\Omega&-1&0&0\\ -x\Omega&0&-1&0\\ 0&0&0&-1\end{pmatrix}.$ (3)
@@
$e^{t}_{0}=e^{x}_{1}=e^{y}_{2}=e^{z}_{3}=1,\quad e^{x}_{0}=y\Omega,\quad e^{y}_{0}=-x\Omega,$ (4)
@@
$\begin{split}\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})&=e^{il\theta}J_{l}(p_{\perp}r){\cal P}_{+},\\ \Phi_{l,-}(p_{\perp};\bm{x}_{\perp})&=e^{i(l+1)\theta}J_{l+1}(p_{\perp}r){\cal P}_{-},\end{split}$ (6)
@@
$\begin{split}\psi(x)&=\int\frac{d^{4}p}{(2\pi)^{4}}e^{-ip^{t}t+ip_{z}z}\sum_{l=-\infty}^{\infty}\Bigl{[}\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})+\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})\Bigr{]}\psi_{l}(p),\\ \psi_{l}(p)&=\int d^{4}xe^{ip^{t}t-ip_{z}z}\Bigl{[}\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp})+\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp})\Bigr{]}\psi(x).\end{split}$ (7)
The modified transformation in Eq. (7) is referred to as the Fourier-Bessel transformation. In the Fourier-Bessel basis, the four-dimensional delta function is expressed as
$\begin{split}\delta^{(4)}(x-x^{\prime})=\int\frac{d^{4}p}{(2\pi)^{4}}e^{-ip^{t}(t-t^{\prime})+ip_{z}(z-z^{\prime})}\sum_{l=-\infty}^{\infty}\left[\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})+\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\right].\end{split}$ (8)
@@
$S(x,x^{\prime})=i\sum_{l=-\infty}^{\infty}\int\frac{d^{4}p}{(2\pi)^{4}}e^{-ip^{t}(t-t^{\prime})+ip_{z}(z-z^{\prime})}\mathcal{S}_{l}(p;\bm{x}_{\perp},\bm{x}_{\perp}^{\prime}),$ (10)
@@
$\begin{split}\mathcal{S}_{l}(p;\bm{x}_{\perp},\bm{x}_{\perp}^{\prime})&=\mathcal{S}_{l,\parallel}(p)\Bigl{[}\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})+\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})\Bigr{]}\\ &\quad+\mathcal{S}_{l,\perp}(p)\Bigl{[}\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})-\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})\Bigr{]},\end{split}$ (11)
$\mathcal{S}_{\parallel}(p)=\frac{\left(p^{t}+\Omega l+\frac{1}{2}\Omega\right)\gamma^{0}-p_{z}\gamma^{3}+m}{\left\{p^{t}+\Omega\left(l+\frac{1}{2}\right)\right\}^{2}-E_{p}^{2}+i\epsilon},\qquad\mathcal{S}_{\perp}(p)=\frac{-ip_{\perp}\gamma^{1}}{\left\{p^{t}+\Omega\left(l+\frac{1}{2}\right)\right\}^{2}-E_{p}^{2}+i\epsilon}.$ (12)
The Dirac propagator obtained above is derived in Ref. [58]. This propagator described with the Fourier-Bessel basis reflects the violation of translational symmetry in various quantities in terms of rotating Dirac fields.
II.2 Definition of susceptibility functions
Before proceeding with the explicit calculation of correlation functions in the rotating frame using the Dirac propagator in Eq. (10), we provide a brief review in this subsection of how susceptibility functions are defined in field theory. To ensure that the susceptibility functions can be defined in any interacting system, we start with the general case of the Lagrangian including an appropriate interaction ${\cal L}_{\rm int}$, as follows:
@@ II.2 Definition of susceptibility functions
$\zeta_{\rm sp}={\zeta_{S}}+i{\zeta_{P}}\gamma_{5},$ (14)
@@
$Z$ $=$ $\int\mathcal{D}\bar{\psi}\mathcal{D}\psi\exp\left[i\int d^{4}x\,\mathcal{L}\right],$ (15)
@@
$\int d^{4}x\,\chi_{c}(x)=(-i)\frac{\partial^{2}\Gamma}{\partial{\zeta_{c}}^{2}}\Biggl{|}_{\zeta_{\rm sp}=0},\qquad c=S,P,$ (17)
@@
$\chi_{c}(x)=\int d^{4}x^{\prime}\langle 0|T{\cal O}_{c}(x){\cal O}_{c}(x^{\prime})|0\rangle,$ (18)
@@
$\begin{split}{\cal O}_{S}=\bar{\psi}\psi,\quad{\cal O}_{P}=\bar{\psi}i\gamma_{5}\psi.\end{split}$ (19)
The susceptibility functions in Eq. (18) represent the quantum fluctuations of the scalar and pseudoscalar fermionic-bilinear fields at zero external momentum.
When deriving the susceptibility in Eq. (17) from the effective action, it always involves a time-space integral. In the usual translationally invariant case, the susceptibility function (18) is constant and thus Eq. (17) is reduced to Eq. (18) multiplied by the four-dimensional volume factor. However, it should be noted that when translational invariance is violated, this reduction is not allowed, requiring a careful treatment of the four-dimensional integral.
II.3 Boundary conditions in finite cylinder

### v2

@@ Abstract
arxiv: 2507.00494
1 Introduction
@@
Rotating QCD matter has garnered significant attention in recent years, particularly in the context of heavy-ion collisions, which create the most vortical fluid known in nature STAR:2017ckg. On the theoretical side, the rigidly rotating system is an idealized setup in order to understand fundamental features of rotating relativistic matter Vilenkin:1978hb; Vilenkin:1979ui; Vilenkin:1980zv, and has already revealed typical rotational responses of QCD matter in terms of the chiral symmetry breaking Chen:2015hfc; Jiang:2016wvv; Ebihara:2016fwa; Chernodub:2016kxh; Wang:2018sur; Chen:2023cjt; Gaspar:2023nqk; Nunes:2024hzy, the deconfinement transition Chernodub:2020qah; Chen:2020ath; Fujimoto:2021xix; Braga:2022yfe; Zhao:2022uxc; Yadav:2022qcl; Chen:2022smf; Chen:2024tkr; Mameda:2023sst; Jiang:2023zzu; Chen:2024edy, both of them Sun:2023kuu; Sun:2024anu; Chen:2024jet; Cao:2023olg, and other possible transitions Huang:2017pqe; Liu:2017spl; Cao:2019ctl; Chen:2019tcp; Zhang:2018ome; see also the corresponding lattice studies in Ref. Chernodub:2022veq; Braguta:2021jgn; Braguta:2022str; Braguta:2023yjn; Braguta:2023iyx; Yang:2023vsw; Wang:2025mmv; Fukushima:2025hmh. Even under such simplification, however, accessible aspects of rotating QCD matter are still limited. In particular, since the first principle simulation under rotation suffers from a sign problem Yamamoto:2013zwa, the sophistication of the model approaches and perturbative computation is an inevitable direction.
@@
A difficulty we face in field-theoretical studies of relativistic rotating matter is the coordinate-dependence of physical observables (see also Refs. Chen:2015hfc; Fukushima:2024tkz for an exceptional case under a strong magnetic field). As demonstrated in preceding works, maintaining causality in a relativistic rotating system necessitates the consideration of a finite-size geometry, typically a cylinder Ambrus:2014uqa. Such a finite-size system violates translational invariance along the radial direction, leading to coordinate-dependent thermodynamic quantities. In more field-theoretical language, propagators of rotating matter are described not by the plane-wave basis, but the Fourier-Bessel basis. As a result, in an interacting theory under rotation, it is necessary to carefully revisit to standard schemes employed in the usual nonrotating case, including the Feynman rules Kuboniwa:2025vpg.
@@
A typical modification is found in effective models describing phase transition under rotation. The standard mean-field analysis is not directly applied to radially inhomogeneous order parameters. One way to bypass this issue is to utilize the local density approximation, where the radial modulation of the order parameter is supposed to be negligible. Under this approximation, for instance, the mean-field analysis for the spontaneous chiral symmetry breaking can be performed parallelly to the usual case, but the constant dynamical mass is replaced with the radial-coordinate dependent one Jiang:2016wvv. A remarkable feature of this analysis is that the self-consistent gap equation involves only the single radial coordinate (through the one-loop of a single propagator), and this is the reason why such a simple analysis works. However, the validity of this strategy becomes subtle when we argue mesonic fluctuations and susceptibilities, which consist of multi-point propagators, and thus involve more radial coordinates than one.
@@
In order to deal with such an issue, Ward-Takahashi (WT) identities would provide a key computational guide. WT identities are highly useful in investigations of QCD properties at finite temperature and density, where Lorentz symmetry is violated by external conditions. For instance, a WT identity in term of the chiral symmetry relates the quark condensate to the two-point correlation function derived from quark bilinear operators corresponding to the pion mode (referred to as the pion susceptibility), and thus plays a crucial role in assessing the consistency within an effective model approach GomezNicola:2019myi; Cui:2021bqf; Cui:2022vsr; Kawaguchi:2023olk. Furthermore, a WT identity for the U(1) axial symmetry associates quark two-point correlation functions with the topological susceptibility, which is originally described by gluonic degrees of freedom GomezNicola:2016ssy; GomezNicola:2017bhm; GomezNicola:2019myi; Kawaguchi:2020qvg; Cui:2021bqf; Cui:2022vsr. Since these WT identities related to the global symmetries is irrelevant to the spacetime geometry, they are expected to be powerful constraints even in the quantum field theory for rotating matter.
@@
The present work aims to analyze fermionic two-point correlation functions of rotating matter in the Fourier-Bessel basis, within the mean-field level of an Nambu-Jona-Lasinio (NJL) model. We compute various quark susceptibilities, including meson, baryon, and topological susceptibilities. In a similar computational manner, we can evaluate the density of moment of inertia in the interacting theory, while the net moment of inertia has already derived in several preceding studies Chernodub:2016kxh; Sun:2023yux; Ahadi:2025rqs. Besides, we show that a WT identity enforces the mesonic susceptibilities to take the similar form to that in the translational invariant systems. These analyses not only establish the groundwork for more comprehensive treatments of quantum fluctuations in rotating matter, but also provides new benchmarks to characterize the rotational response to QCD phase transition. Indeed, without rotation, the phase structure of hot QCD has been explored, revealing the chiral crossover where the quark condensate decreases smoothly with increasing temperature Aoki:2006we; Bhattacharya:2014ara. Additionally, meson susceptibilities, derived from quark bilinear correlation functions, have also been investigated and have provided valuable insights into meson property fluctuations and symmetry breaking, including chiral $SU(2)$ and $U(1)$ axial symmetries HotQCD:2012vvd; Bhattacharya:2014ara; Buchoff:2013nra; Aoki:2020noz; Aoki:2021qws. Given this fact, our study extends the understanding of rotating QCD matter by elucidating how rotation modifies these susceptibility functions and the associated symmetry restoration patterns.
@@
This paper is organized as follows. In section 2, we present the theoretical framework in the rotating cylindrical frame. We briefly review the formulation of the Dirac propagator described in the Fourier-Bessel basis and introduce the definition of the susceptibility functions. The boundary condition in a finite cylinder is also provided. In section 3, we analyze the susceptibility functions within the NJL model under the mean-field approximation. We examine the implementation of the local density approximation in the current framework and explore the associated constraints on the quark propagator. This approximation simplifies the calculation of the susceptibility functions and, guided by a WT identity, makes it possible to evaluate meson susceptibilities. In section 4, we show our numerical results on the susceptibility functions at finite temperature under rotation. Finally, we summarize our findings and discuss possible future directions in section 5.
2 Theoretical setup
@@
2.1 Dirac propagator in the Fourier-Bessel basis
@@
$\Bigl[i\gamma^{\mu}(\partial_{\mu}+\Gamma_{\mu})-m\Bigr]\psi=0,$ (2.1)
@@
$g_{\mu\nu}=\eta_{ij}e^{i}_{\mu}e^{j}_{\nu}.$ (2.2)
In this paper, we consider a system rotating around the $z$-axis with an angular velocity, $^{\bm{\Omega}}=\Omega\hat{\bm{z}}$, and thus employ the corresponding metric as
$g_{\mu\nu}=\begin{pmatrix}1-(x^{2}+y^{2})\Omega^{2}&y\Omega&-x\Omega&0\\ y\Omega&-1&0&0\\ -x\Omega&0&-1&0\\ 0&0&0&-1\end{pmatrix}$. (2.3)
@@
$e^{t}_{0}=e^{x}_{1}=e^{y}_{2}=e^{z}_{3}=1,\quad e^{x}_{0}=y\Omega,\quad e^{y}_{0}=-x\Omega,$ (2.4)
@@
$\begin{split}\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})&=e^{il\theta}J_{l}(p_{\perp}r){\cal P}_{+},\\ \Phi_{l,-}(p_{\perp};\bm{x}_{\perp})&=e^{i(l+1)\theta}J_{l+1}(p_{\perp}r){\cal P}_{-},\end{split}$ (2.6)
@@
$\begin{split}\psi(x)&=\int\frac{d^{4}p}{(2\pi)^{4}}e^{-ip^{t}t+ip_{z}z}\sum_{l=-\infty}^{\infty}\Bigl[\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})+\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})\Bigr]\psi_{l}(p),\\ \psi_{l}(p)&=\int d^{4}xe^{ip^{t}t-ip_{z}z}\Bigl[\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp})+\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp})\Bigr]\psi(x).\end{split}$ (2.7)
The modified transformation in Eq. (2.7) is referred to as the Fourier-Bessel transformation. In the Fourier-Bessel basis, the four-dimensional delta function is expressed as
$\begin{split}\delta^{(4)}(x-x^{\prime})&=\int\frac{d^{4}p}{(2\pi)^{4}}e^{-ip^{t}(t-t^{\prime})+ip_{z}(z-z^{\prime})}\\ &\quad\times\sum_{l=-\infty}^{\infty}\left[\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})+\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\right].\end{split}$ (2.8)
@@
$S(x,x^{\prime})=i\sum_{l=-\infty}^{\infty}\int\frac{d^{4}p}{(2\pi)^{4}}e^{-ip^{t}(t-t^{\prime})+ip_{z}(z-z^{\prime})}\mathcal{S}_{l}(p;\bm{x}_{\perp},\bm{x}_{\perp}^{\prime}),$ (2.10)
@@
$\begin{split}\mathcal{S}_{l}(p;\bm{x}_{\perp},\bm{x}_{\perp}^{\prime})&=\mathcal{S}_{l,\parallel}(p)\Bigl[\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})+\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})\Bigr]\\ &\quad+\mathcal{S}_{l,\perp}(p)\Bigl[\Phi_{l,+}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,-}(p_{\perp};\bm{x}_{\perp})-\Phi_{l,-}^{\dagger}(p_{\perp};\bm{x}_{\perp}^{\prime})\Phi_{l,+}(p_{\perp};\bm{x}_{\perp})\Bigr],\end{split}$ (2.11)
$\mathcal{S}_{\parallel}(p)=\frac{\left(p^{t}+\Omega l+\frac{1}{2}\Omega\right)\gamma^{0}-p_{z}\gamma^{3}+m}{\left\{p^{t}+\Omega\left(l+\frac{1}{2}\right)\right\}^{2}-E_{p}^{2}+i\epsilon},\qquad\mathcal{S}_{\perp}(p)=\frac{-ip_{\perp}\gamma^{1}}{\left\{p^{t}+\Omega\left(l+\frac{1}{2}\right)\right\}^{2}-E_{p}^{2}+i\epsilon}.$ (2.12)
The Dirac propagator obtained above is derived in Ref. Ambrus:2015lfr. This propagator described with the Fourier-Bessel basis reflects the violation of translational symmetry in various quantities in terms of rotating Dirac fields.
2.2 Definition of susceptibility functions
Before proceeding with the explicit calculation of correlation functions in the rotating frame using the Dirac propagator in Eq. (2.10), we provide a brief review in this subsection of how susceptibility functions are defined in field theory. To ensure that the susceptibility functions can be defined in any interacting system, we start with the general case of the Lagrangian including an appropriate interaction ${\cal L}_{\rm int}$, as follows:
@@ 2.2 Definition of susceptibility functions
$\zeta_{\rm sp}={\zeta_{S}}+i{\zeta_{P}}\gamma_{5},$ (2.14)
@@
$Z$ $=$ $\int\mathcal{D}\bar{\psi}\mathcal{D}\psi\exp\left[i\int d^{4}x\,\mathcal{L}\right],$ (2.15)
@@
$\int d^{4}x\,\chi_{c}(x)=(-i)\frac{\partial^{2}\Gamma}{\partial{\zeta_{c}}^{2}}\Biggl|_{\zeta_{\rm sp}=0},\qquad c=S,P,$ (2.17)
@@
$\chi_{c}(x)=\int d^{4}x^{\prime}\langle 0|T{\cal O}_{c}(x){\cal O}_{c}(x^{\prime})|0\rangle,$ (2.18)
@@
$\begin{split}{\cal O}_{S}=\bar{\psi}\psi,\quad{\cal O}_{P}=\bar{\psi}i\gamma_{5}\psi.\end{split}$ (2.19)
The susceptibility functions in Eq. (2.18) represent the quantum fluctuations of the scalar and pseudoscalar fermionic-bilinear fields at zero external momentum.
When deriving the susceptibility in Eq. (2.17) from the effective action, it always involves a time-space integral. In the usual translationally invariant case, the susceptibility function (2.18) is constant and thus Eq. (2.17) is reduced to Eq. (2.18) multiplied by the four-dimensional volume factor. However, it should be noted that when translational invariance is violated, this reduction is not allowed, requiring a careful treatment of the four-dimensional integral.
2.3 Boundary conditions in finite cylinder